Are all employees equally proven to having accidents? To investigate this hypothesis, Parry (1985) looked at a light manufacturing plant and classified the accidents by type and by age of the employee.

AgeAccident (sprain)Accident (burn)Accident (cut)    Under 25917525 and over611312

 



1
Expert's answer
2021-08-09T11:23:18-0400

Whole problem



It is needed to test, whether the all employees equally prone to having accidents.

The null and alternative hypotheses for the test are:

H0: Accident type seem to be independent of age.

H1: Accident type does not seem to be independent of age.

The expected frequency formula for a particular cell:

Expected frequency "= \frac{Row \; total \times Column \; total}{Overall \; frequency}"

Consider Oi is the observed frequency for ith cell and Ei is the expected frequency for ith cell. There are total 2 rows (r) and 3 columns (c).

The chi-square test:

"χ^2 = \sum \frac{(O_i -E_i)^2}{E_i}"

Degree of freedom (r-1)(c-1)

"χ^2 = 20.78 \\

df = (2-1)(3-1)=2"

Consider that the level of significance is α = 0.05.

Using the Excel formula =CHISQ.DIST.RT(20.78,2) the P-value is approximately 0.

Decision rule:

If P-value is less than or equal to the level of significance, then reject the null hypothesis. Otherwise, fail to reject the null hypothesis.

Since the P-value of 0 is less than the level of significance of 0.05, reject H0.

Thus, there is enough evidence to conclude that accident type does not seem to be independent of age, at 5% level of significance.

Correct option is C.